Suppose Y is a random variable with pdf f(y)=ay²-b,0<y<3

Find a and b given that E(Y)=15/7


1
Expert's answer
2021-05-07T09:29:06-0400
"\displaystyle\int_{0}^3(ay^2-b)=1"

"[\dfrac{ay^3}{3}-by]\begin{matrix}
 3 \\
 0
\end{matrix}=1"

"9a-3b=1"

"E(Y)=\displaystyle\int_{0}^3y(ay^2-b)=[\dfrac{ay^4}{4}-\dfrac{by^2}{2}]\begin{matrix}
 3 \\
 0
\end{matrix}"

"=\dfrac{81a}{4}-\dfrac{9b}{2}=\dfrac{15}{7}"

"b=3a-\dfrac{1}{3}"

"27a-18a+2=\dfrac{20}{7}"

"9a=\dfrac{6}{7}"

"a=\dfrac{2}{21}"


"b=-\dfrac{1}{21}"